# natural language processing; greenwashing detection; corporate climate-related greenwashing; climate pretrained models; climate topic detection; green claim detection; climate QA; climate-related deceptive techniques; enviromental performance prediction copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing

### v1

@@ 4.5. Green Claim Detection
Green Claims (Woloszyn et al., 2021) Marketing Tweets Environmental (or green) advertisements refer to all appeals that include ecological, environmental sustainability, or nature-friendly messages that target the needs and desires of environmentally concerned stakeholders.
Environmental Claims (Stammbach et al., 2023) Paragraph from reports Environmental claims refer to the practice of suggesting or otherwise creating the impression […] that a product or a service is environmentally friendly (i.e., it has a positive impact on the environment) or is less damaging to the environment than competing goods or services […] In our case, claims relate to products, services, or specific corporate environmental performance.
Table 5. Label definitions of the datasets on green claims detection.
@@
The most notable approaches aiming at identifying climate-related claims introduced annotated datasets (Woloszyn et al., 2021; Stammbach et al., 2023). Woloszyn et al. (2021) proposed to focus on social media marketing through the detection of green claims in tweets. Stammbach et al. (2023) proposed a dataset for environmental claim detection in paragraphs from corporate reports.
Supervised Solutions
Both Stammbach et al. (2023) and Woloszyn et al. (2021) evaluated fine-tuned Transformer models on the datasets. Stammbach et al. (2023) also evaluated classical approaches and experimented with a model fine-tuned on general claim detection (Arslan et al., 2020) applied to their domain-specific task (see Table LABEL:tab:models in the Appendix for details). Woloszyn et al. (2021) and Stammbach et al. (2023) demonstrated good performances with fine-tuned Transformer (above 84%), yet they have very different Inter-Annotator Agreement (IAA), with Krippendorff’s $\alpha=0.8223$ for Woloszyn et al. (2021) and $\alpha=0.47$ for Stammbach et al. (2023). This might be explained by the difference in context: marketing tweets might be easier to understand for humans and more self-sufficient compared to paragraphs extracted from reports. Additionally, Woloszyn et al. (2021) experimented with adversarial attack, showing the sensibility to character-swap and word-swap. This shows that the models tend to rely heavily on particular words, showing only a superficial understanding. While it is difficult to generalize that conclusion given the small size of the dataset, it still indicates that BERT-like architecture can over-fit or rely on superficial cues instead of building an accurate representation of the paragraph. It is essential to assess model performance in challenging scenarios, where the presence of nonsensical or noisy inputs can reveal the fragility of model comprehension.
@@
Green claims have been analyzed in company reporting (Stammbach et al., 2023) and social media communication (Woloszyn et al., 2021). Fine-tuned models can solve the task rather well, even if challenges remain:
  • Inter annotator agreement: The annotator agreement remains low for green claim detection in reports (Stammbach et al., 2023).
  • Integration with existing claim detection literature: While claim detection is a well-established field with extensive literature, studies on environmental claims do not fully connect with existing research on claim detection. In particular, they do not distinguish between Claims, Verifiable Claims (claims that can be checked), and Check-worthy Claims (claims that are interesting to verify) (Panchendrarajan and Zubiaga, 2024).
@@
4.6. Green Claim Characteristics
@@
Implicit/Explicit Green Claims (Woloszyn et al., 2021) Marketing Tweets Implicit green claims raise the same ecological and environmental concerns as explicit green claims (see definition in Section 4.5), but without showing any commitment from the company. If the tweet does not contain a green claim then No Claim.
Specificity (Bingler et al., 2023) Paragraph from reports A paragraph is Specific if it includes clear, tangible, and firm-specific details about events, goals, actions, or explanations that directly impact or clarify the firm’s operations, strategy, or objectives. Non-specific otherwise.
Commitments & Actions(Bingler et al., 2023) Paragraph from reports A paragraph is a commitment or an action if it contains targets for the future or actions already taken in the past.
Net Zero/Reduction (Schimanski et al., 2023a) Paragraph from Net Zero Tracker (Lang et al., 2023) The paragraph contains either a Net-Zero target, a Reduction target or no target (None)
Table 6. Label definitions the datasets related to characterization of green claims.
@@
Existing works characterized both the content and the form of claims. On the content dimension, there are datasets for identifying if the claim is about an action or a commitment (Bingler et al., 2023; Schimanski et al., 2023a). Bingler et al. (2023) introduce a dataset for the identification of commitments and actions (Yes/No), and Schimanski et al. (2023a) released one for the identification of reduction targets (net zero, reduction, general). On the form dimension, Woloszyn et al. (2021) characterized claims as implicit or explicit, and Bingler et al. (2023) annotated their dataset on the specificity of claims (Specific/Not specific). Bingler et al. (2023) ultimately used the specificity and commitment/action characteristics to identify cheap talks related to climate disclosure.
Supervised Solutions
Woloszyn et al. (2021); Schimanski et al. (2023a) and Bingler et al. (2023) experimented with fine-tuned Transformers. Bingler et al. (2023) also evaluated classical baselines (SVM, NB), and Schimanski et al. (2023a) experimented with GPT-3.5-turbo (see Table LABEL:tab:models in the Appendix for details). All approaches reached good performances, in particular fine-tuned Transformers reaching performances above 80%. The exception remains the Specificity classification with a F1-score of 77% (close to the baselines). This low performance might be intrinsic to the task. Indeed, humans have disagreement when distinguishing specific and unspecific claims: Bingler et al. (2023) measured a low IAA on Specificity (Krippendorff’s $\alpha=0.1703$). This might indicate that the task is not well defined in the first place.
@@
The high performances of fine-tuned models indicate that this type of task is solved. Furthermore, the demonstrated ability of GPT-3.5 as a zero-shot classifier (Schimanski et al., 2023a) highlights the potential of large language models to classify these characteristics without requiring extensive annotated datasets. These characteristics are particularly useful for identifying statements that may indicate greenwashing. For example, Bingler et al. (2023) utilized the proportion of non-specific commitments as a Cheap Talk Index demonstrating their practical applications. However, characteristics such as specificity are inherently ambiguous and subjective, as evidenced by the low inter-annotator agreement (IAA). Therefore, future research could focus on disambiguating what constitutes a specific statement from a non-specific one more objectively.
4.7. Green Stance Detection
@@
ClimateFEVER (evidence)  (Leippold and Diggelmann, 2020) A claim and an evidence sentence from Wikipedia Support, Refutes, Not Enough Information Determines the relation between a claim and a single evidence sentence
LobbyMap (Stance) (Morio and Manning, 2023) Page from a company communications (report, press release, …) Strongly supporting, Supporting, No or mixed position, Not supporting, Opposing Given the policy and the page, classifies the stance
Global Warming Stance Detection (GWSD) (Luo et al., 2020) Sentences from news about global warming Stance of the evidence (Agree, Disagree, Neutral) toward the claim: Climate-Change is a serious concern.
ClimateStance (Vaid et al., 2022) Tweets posted during COP25 filtered by keywords (relevant to climate-change) Stance towards climate change prevention: Favor, Against, Ambiguous. (Stance used as a broad notion including sentiment, evaluation, appraisal, …)
Stance on Remediation Effort (Lai and Chen, 2023) Texts extracted from the TCFD sections of the financial reports The text indicates support for climate change remediation efforts or refutation of such efforts. (No guidelines)
@@
ClimateFEVER (claim)  (Leippold and Diggelmann, 2020) A claim and multiple evidence sentences from Wikipedia Support, Refutes, Debated, Not Enough Information Determines if a claim is supported by a set of retrieved evidence sentences
LobbyMap (Page) (Morio and Manning, 2023) Page from a company communications (report, press release, …) 1/0 Contains a stance on a remediation policy
LobbyMap (Query) (Morio and Manning, 2023) GHG emission regulation, Renewable energy, Carbon tax, … Classifies the remediation policy
Table 7. Summary of datasets, labels, scopes, and sources for climate-related stance detection.
@@
The first subtask is collecting the evidences when the claims are already available (e.g. if they were collected manually). In order to build these datasets, research can rely on simple heuristics such as downloading all tweets published during the COP25 filtered with keywords (Vaid et al., 2022). However, other researchers performed a more elaborate procedure. Leippold and Diggelmann (2020) proposed a pipeline for collecting evidence. To retrieve relevant evidence from Wikipedia for a given claim, the pipeline involves three steps: document-level retrieval using entity-linking and BM25 to identify top articles, sentence-level retrieval using sentence embeddings trained on the FEVER dataset to extract relevant sentences, and sentence re-ranking using a pretrained ALBERT model to classify and rank evidence based on relevance. The second one encompasses the broader process: identifying claims and finding evidence, before predicting the relation. Wang et al. (2021) used a generic claim detection model trained on ClaimBuster (Arslan et al., 2020) for the claims and Google Search API to collect evidence. Finally, the last subtask is to train multiple models on each step: identifying evidences, identifying the claim and classifying the stance. This is the approach used by (Morio and Manning, 2023).
@@
As described previously, researchers built datasets of evidence related to one claim while focusing stance in tweets on seriousness of climate change such as Global Warming Stance Detection (GWSD) (Luo et al., 2020), ClimateStance (Vaid et al., 2022) stance on climate change of tweets posted during the COP25, or Lai and Chen (2023)’s dataset which focuses on the stance toward climate change remediation efforts. Another research initiative built a dataset of claim-evidence pairs, ClimateFEVER, and then trained a model on the stance classification (Leippold and Diggelmann, 2020; Wang et al., 2021). Finally, Morio and Manning (2023) proposed a dataset to assess corporate policy engagement built upon LobbyMap, which tackles the 3 steps: finding pages with evidence, identifying the claims targeted by that evidence, and classifying the stance.
Supervised Solutions
The solution proposed for to tackle stance classification are fine-tuned Transformer models (Vaid et al., 2022; Vaghefi et al., 2022; Xiang and Fujii, 2023; Morio and Manning, 2023; Webersinke et al., 2022; Wang et al., 2021; Spokoyny et al., 2023; Lai and Chen, 2023; Luo et al., 2020) and classical approaches (Spokoyny et al., 2023; Morio and Manning, 2023; Luo et al., 2020). Vaid et al. (2022) also experimented with FastText and Xiang and Fujii (2023) with LSTM-based models (See Table LABEL:tab:models in the Appendix for details). The performances are quite heterogeneous, Lai and Chen (2023) reaching F1-scores around 90% on classification of stance on remediation efforts, Luo et al. (2020) around 72% on GWSD, while performance on ClimateStance could not exceed 60%. Performances on ClimateFEVER are also quite low, however, when selecting only non-ambiguous examples, Xiang and Fujii (2023) could reach performances around 80%. Finally, performances on LobbyMap (Morio and Manning, 2023) are quite low (between 31% and 57.3% depending on the strictness of the metric). Overall, the performances show that the datasets are challenging (see Table 19 in the Appendix for details).
@@
Datasets such as ClimateFEVER (Leippold and Diggelmann, 2020) and LobbyMap (Morio and Manning, 2023) rely on automated construction methods that may not ensure exhaustivity of evidence coverage. ClimateFEVER utilizes BM25 and Wikipedia for evidence retrieval, which could result in missing relevant information, particularly when compared to more advanced retrieval methods. LobbyMap relies on the LobbyMap.org website, which was not designed to be exhaustive. These limitations should be further investigated.
@@
4.8. Question Answering
Question answering (QA) is a known task:
@@ 4.8. Question Answering
ClimaQA  (Spokoyny et al., 2023) The text from a response to one of the CDP questions; and one of the questions from the CDP questionnaire 1: the response answers this question 0: The response does not answer this question, but another one
ClimateQA (Luccioni et al., 2020) Sentence from reports and a question based on the TCFD recommendations 1: the sentence answers the question 0: The sentences does not answer the question
ClimaINS  (Spokoyny et al., 2023) The text from a response to one of the questions from the NAIC questionnaire MANAGE, RISK PLAN, MITIGATE, ENGAGE, ASSESS, RISKS The labels correspond the 8 questions asked in the NAIC questionnaires
Table 8. Summary of datasets, labels, scopes, and sources for climate-related QA datasets.
QA can be used for climate-specific applications such as structuring the information related to climate change from documents (Luccioni et al., 2020; Schimanski et al., 2023a), building chatbots with climate-related knowledge to make it more accessible to non-experts (Vaghefi et al., 2023; Rony et al., 2022), or helping identify potentially misleading information (Ni et al., 2023).
@@
The QA process can be divided into two main steps (Kraus et al., 2023): the retrieval step and the answer generation step. The retrieval step involves locating the relevant information or answer to the user’s query in external documents. The answer generation step involves formulating a response to the user’s query from (1) the information retrieved from documents in the first step or (2) the pre-trained internal knowledge of a model.
@@
Retriever systems can be generic, using techniques such as sentence similarity, BM25, or document tags to filter documents and retrieve specific passages. Schimanski et al. (2024b) published ClimRetrieve, a benchmark for climate-related information retrieval in corporate reports. They found that simple embedding approaches are limited. Therefore, several works have specifically focused on improving the identification of answers within climate-related documents (Luccioni et al., 2020; Spokoyny et al., 2023). Luccioni et al. (2020) reformulated TCFD recommendations as questions and annotated reports to identify answers to the questions. The dataset includes questions paired with potential answers, each labeled to indicate whether the answer addresses the question or not. Based on the QA model trained on this dataset, they also proposed a tool called ClimateQA. Spokoyny et al. (2023) focused on questions from the CDP questionnaire and the NAIC Climate Risk Disclosure survey. As the responses to those questionnaires are publicly available, they constructed two datasets, ClimaINS and ClimaQA. ClimaQA is similar to ClimateQA (question/potential answer pairs). On the contrary, ClimaINS is a QA dataset framed as a classification task: it contains the responses from the NAIC survey, and each response is labeled as answering one of the 8 questions of the survey. Once a model has been trained on these datasets, it can be used to search for answers to the questions in other documents (effectively retrieving the information from an unstructured document). For ClimaINS, the authors experimented with multiple fine-tuned Transformers. As ClimaQA is framed as a retrieval task, the authors used BM25, sentence-BERT, and ClimateBERT to rank possible answers. Using the methodology evaluated on ClimaQA, they proposed a system to extract responses to CDP questions directly from unstructured reports and automatically fill the questionnaire (Spokoyny et al., 2023). Luccioni et al. (2020) trained and evaluated a RoBERTa-based model called ClimateQA on finding answers to TCFD-based questions. They all reached good performance but with room for improvement (see Table 21 in the Appendix for details), demonstrating the feasibility of the task but also the need for further research.
@@
After the retrieval, the second step of question answering is to generate an answer for the question (given input resources or not). The generation can be simple such as finding a particular information in a paragraph. For example, Schimanski et al. (2023a) used a QA model (RoBERTa SQuaD v2 (Rajpurkar et al., 2016)) to extract the target year, the percentage of reduction, and the baseline year from the reduction target of a company from an input text that contains an emission reduction target. While it is possible to rely on generalist models, as they are proficient few-shot learners (Brown et al., 2020; Thulke et al., 2024), multiple studies proposed using domain-specific models. Climate Bot (Rony et al., 2022) built a dataset of climate-specific QA. Given a question and a document (scientific/news), the model should find the span answering the question. Mullappilly et al. (2023) proposed a dataset of question-answer pairs based on ClimaBench (Spokoyny et al., 2023) and CCMRC (Rony et al., 2022) to train a model to generate answers. Thulke et al. (2024) introduced a climate-specific corpus of prompt/completion pairs for Instruction Fine-Tuning created by experts and non-experts. (Rony et al., 2022) finetuned an ALBERT model on their climate-specific QA dataset. Mullappilly et al. (2023) trained a Vicuna-7B on their climate-specific answer generation dataset. And Thulke et al. (2024) trained LLama-2-based models (ClimateGPT) on their domain-specific prompt/completion dataset. They also evaluated their ClimateGPT model alongside multiple LLMs on climate-related benchmarks. They found that the domain-specific models outperformed generalist ones (see Table 21 in the Appendix for details).
@@
By combining the last two components (a retriever and a QA system), one can develop Retrieval-Augmented Generation (RAG) systems. Rony et al. (2022) provide the first example of a RAG system in climate-related tasks based on sentence-BERT to retrieve documents and AlBERT to identify the answer. More recently, Ni et al. (2023) proposed ChatReport, a methodology based on ChatGPT for analyzing corporate reports through the lens of TCFD questions using RAG. The reports, the answers to the TCFD questions, and a TCFD conformity assessment are chunked and stored in a vector store for easy retrieval. Each chunk and/or answer can be retrieved and injected into the prompt. They added in the prompt the notions of greenwashing and cheap talk to invite the model to provide a critical analysis of the retrieved answers. Similarly, Vaghefi et al. (2023) introduced ChatClimate, a RAG-based pipeline to augment GPT-4 with knowledge about IPCC reports. While the previous LLM-based approaches rely on closed-source models, Mullappilly et al. (2023) and Thulke et al. (2024) proposed RAG-based pipelines relying on open-source models. Rony et al. (2022) is the only study that evaluated their RAG pipeline using classical metrics (F1-score, BLEU, METEOR). Unfortunately, those metrics are not well aligned with human judgment for text generation (Chen et al., 2019). Therefore Vaghefi et al. (2023) and Mullappilly et al. (2023) relied on human or ChatGPT evaluations. They conclude that their approaches improve on non-domain-specific RAG systems.

### v2

@@ 4.5 Green Claim Detection
Green Claims vinicius_woloszyn_towards_2021 Marketing Tweets Environmental (or green) advertisements refer to all appeals that include ecological, environmental sustainability, or nature-friendly messages that target the needs and desires of environmentally concerned stakeholders.
Environmental Claims stammbach_environmental_2023 Paragraph from reports Environmental claims refer to the practice of suggesting or otherwise creating the impression […] that a product or a service is environmentally friendly (i.e., it has a positive impact on the environment) or is less damaging to the environment than competing goods or services […] In our case, claims relate to products, services, or specific corporate environmental performance.
@@
The most notable approaches aiming at identifying climate-related claims introduced annotated datasets vinicius_woloszyn_towards_2021, stammbach_environmental_2023. vinicius_woloszyn_towards_2021 proposed to focus on social media marketing through the detection of green claims in tweets. stammbach_environmental_2023 proposed a dataset for environmental claim detection in paragraphs from corporate reports.
Models
Both stammbach_environmental_2023 and vinicius_woloszyn_towards_2021 evaluated fine-tuned Transformer models on the datasets. stammbach_environmental_2023 also evaluated classical approaches and experimented with a model fine-tuned on general claim detection Arslan_Hassan_Li_Tremayne_2020 applied to their domain-specific task. vinicius_woloszyn_towards_2021 and stammbach_environmental_2023 demonstrated good performances with fine-tuned Transformer (above 84%), yet they have very different Inter-Annotator Agreement (IAA), with Krippendorff’s $\alpha=0.8223$ for vinicius_woloszyn_towards_2021 and $\alpha=0.47$ for stammbach_environmental_2023. This might be explained by the difference in context: marketing tweets might be easier to understand for humans and more self-sufficient compared to paragraphs extracted from reports. Additionally, vinicius_woloszyn_towards_2021 experimented with adversarial attack, showing the sensibility to character-swap and word-swap. This shows that the models tend to rely heavily on particular words, showing only a superficial understanding. While it is difficult to generalize that conclusion given the small size of the dataset, it still indicates that BERT-like architecture can over-fit or rely on superficial cues instead of building an accurate representation of the paragraph. It is essential to assess model performance in challenging scenarios, where the presence of nonsensical or noisy inputs can reveal the fragility of model comprehension.
@@
Green claims have been analyzed in company reporting [stammbach_environmental_2023] and social media communication [vinicius_woloszyn_towards_2021]. Fine-tuned models can solve the task rather well, even if challenges remain:
  • Inter annotator agreement: The annotator agreement remains low for green claim detection in reports [stammbach_environmental_2023].
  • Integration with existing claim detection literature: While claim detection is a well-established field with extensive literature, studies on environmental claims do not fully connect with existing research on claim detection. In particular, they do not distinguish between Claims, Verifiable Claims (claims that can be checked), and Check-worthy Claims (claims that are interesting to verify) [panchendrarajan2024claim].
@@
4.6 Green Claim Characteristics
@@
Table 8: Label definitions the datasets related to characterization of green claims.
@@
Implicit/Explicit Green Claims vinicius_woloszyn_towards_2021 Marketing Tweets Implicit green claims raise the same ecological and environmental concerns as explicit green claims (see definition in Section 4.5), but without showing any commitment from the company. If the tweet does not contain a green claim then No Claim.
Specificity bingler2023cheaptalkspecificitysentiment Paragraph from reports A paragraph is Specific if it includes clear, tangible, and firm-specific details about events, goals, actions, or explanations that directly impact or clarify the firm’s operations, strategy, or objectives. Non-specific otherwise.
Commitments & Actionsbingler2023cheaptalkspecificitysentiment Paragraph from reports A paragraph is a commitment or an action if it contains targets for the future or actions already taken in the past.
Net Zero/Reduction tobias_schimanski_climatebert-netzero_2023 Paragraph from Net Zero Tracker [netzerotracker2023] The paragraph contains either a Net-Zero target, a Reduction target or no target (None)
@@
Existing works characterized both the content and the form of claims. On the content dimension, there are datasets for identifying if the claim is about an action or a commitment [bingler2023cheaptalkspecificitysentiment, tobias_schimanski_climatebert-netzero_2023]. bingler2023cheaptalkspecificitysentiment introduce a dataset for the identification of commitments and actions (Yes/No), and tobias_schimanski_climatebert-netzero_2023 released one for the identification of reduction targets (net zero, reduction, general). On the form dimension, vinicius_woloszyn_towards_2021 characterized claims as implicit or explicit, and bingler2023cheaptalkspecificitysentiment annotated their dataset on the specificity of claims (Specific/Not specific). bingler2023cheaptalkspecificitysentiment ultimately used the specificity and commitment/action characteristics to identify cheap talks related to climate disclosure.
Models
vinicius_woloszyn_towards_2021, tobias_schimanski_climatebert-netzero_2023 and bingler2023cheaptalkspecificitysentiment experimented with fine-tuned Transformers. bingler2023cheaptalkspecificitysentiment also evaluated classical baselines (SVM, NB), and tobias_schimanski_climatebert-netzero_2023 experimented with GPT-3.5-turbo. All approaches reached good performances, in particular fine-tuned Transformers reaching performances above 80%. The exception remains the Specificity classification with a F1-score of 77% (close to the baselines). This low performance might be intrinsic to the task. Indeed, humans have disagreement when distinguishing specific and unspecific claims: bingler2023cheaptalkspecificitysentiment measured a low IAA on Specificity (Krippendorff’s $\alpha=0.1703$). This might indicate that the task is not well defined in the first place.
@@
The high performances of fine-tuned models indicate that this type of task is solved. Furthermore, the demonstrated ability of GPT-3.5 as a zero-shot classifier tobias_schimanski_climatebert-netzero_2023 highlights the potential of large language models to classify these characteristics without requiring extensive annotated datasets. These characteristics are particularly useful for identifying statements that may indicate greenwashing. For example, bingler2023cheaptalkspecificitysentiment utilized the proportion of non-specific commitments as a Cheap Talk Index demonstrating their practical applications. However, characteristics such as specificity are inherently ambiguous and subjective, as evidenced by the low inter-annotator agreement (IAA). Therefore, future research could focus on disambiguating what constitutes a specific statement from a non-specific one more objectively.
4.7 Green Stance Detection
@@
Table 9: Summary of datasets, labels, scopes, and sources for climate-related stance detection.
@@
ClimateFEVER (evidence) [diggelmann_climate-fever_2020] A claim and an evidence sentence from Wikipedia Support, Refutes, Not Enough Information Determines the relation between a claim and a single evidence sentence
LobbyMap (Stance) [morio2023an] Page from a company communications (report, press release, …) Strongly supporting, Supporting, No or mixed position, Not supporting, Opposing Given the policy and the page, classifies the stance
Global Warming Stance Detection (GWSD) [luo_detecting_2020] Sentences from news about global warming Stance of the evidence (Agree, Disagree, Neutral) toward the claim: Climate-Change is a serious concern.
ClimateStance [vaid-etal-2022-towards] Tweets posted during COP25 filtered by keywords (relevant to climate-change) Stance towards climate change prevention: Favor, Against, Ambiguous. (Stance used as a broad notion including sentiment, evaluation, appraisal, …)
Stance on Remediation Effort [lai_using_2023] Texts extracted from the TCFD sections of the financial reports The text indicates support for climate change remediation efforts or refutation of such efforts. (No guidelines)
@@
ClimateFEVER (claim) [diggelmann_climate-fever_2020] A claim and multiple evidence sentences from Wikipedia Support, Refutes, Debated, Not Enough Information Determines if a claim is supported by a set of retrieved evidence sentences
LobbyMap (Page) [morio2023an] Page from a company communication (report, press release, etc) Binary Contains a stance on a remediation policy
LobbyMap (Query) [morio2023an] GHG emission regulation, Renewable energy, Carbon tax, etc) Classifies the remediation policy
@@
The first subtask is collecting the evidences when the claims are already available (e.g. if they were collected manually). In order to build these datasets, research can rely on simple heuristics such as downloading all tweets published during the COP25 filtered with keywords vaid-etal-2022-towards. However, other researchers performed a more elaborate procedure. diggelmann_climate-fever_2020 proposed a pipeline for collecting evidence. To retrieve relevant evidence from Wikipedia for a given claim, the pipeline involves three steps: document-level retrieval using entity-linking and BM25 to identify top articles, sentence-level retrieval using sentence embeddings trained on the FEVER dataset to extract relevant sentences, and sentence re-ranking using a pretrained ALBERT model to classify and rank evidence based on relevance. The second one encompasses the broader process: identifying claims and finding evidence, before predicting the relation. Wang2021EvidenceBA used a generic claim detection model trained on ClaimBuster Arslan_Hassan_Li_Tremayne_2020 for the claims and Google Search API to collect evidence. Finally, the last subtask is to train multiple models on each step: identifying evidences, identifying the claim and classifying the stance. This is the approach used by morio2023an.
@@
As described previously, researchers built datasets of evidence related to one claim while focusing stance in tweets on seriousness of climate change such as Global Warming Stance Detection (GWSD) luo_detecting_2020, ClimateStance vaid-etal-2022-towards stance on climate change of tweets posted during the COP25, or lai_using_2023’s dataset which focuses on the stance toward climate change remediation efforts. Another research initiative built a dataset of claim-evidence pairs, ClimateFEVER, and then trained a model on the stance classification diggelmann_climate-fever_2020, Wang2021EvidenceBA. Finally, morio2023an proposed a dataset to assess corporate policy engagement built upon LobbyMap, which tackles the 3 steps: finding pages with evidence, identifying the claims targeted by that evidence, and classifying the stance.
Models
The solution proposed for to tackle stance classification are fine-tuned Transformer models vaid-etal-2022-towards, vaghefi2022deep, xiang_dare_2023, morio2023an, nicolas_webersinke_climatebert_2021, Wang2021EvidenceBA, spokoyny2023answering, lai_using_2023, luo_detecting_2020 and classical approaches spokoyny2023answering, morio2023an, luo_detecting_2020. vaid-etal-2022-towards also experimented with FastText and xiang_dare_2023 with LSTM-based models . The performances are quite heterogeneous, lai_using_2023 reaching F1-scores around 90% on classification of stance on remediation efforts, luo_detecting_2020 around 72% on GWSD, while performance on ClimateStance could not exceed 60%. Performances on ClimateFEVER are also quite low, however, when selecting only non-ambiguous examples, xiang_dare_2023 could reach performances around 80%. Finally, performances on LobbyMap [morio2023an] are quite low (between 31% and 57.3% depending on the strictness of the metric). Overall, the performances show that the datasets are challenging.666see Table 20 in the Appendix for details.
@@
Datasets such as ClimateFEVER [diggelmann_climate-fever_2020] and LobbyMap [morio2023an] rely on automated construction methods that may not ensure exhaustivity of evidence coverage. ClimateFEVER utilizes BM25 and Wikipedia for evidence retrieval, which could result in missing relevant information, particularly when compared to more advanced retrieval methods. LobbyMap relies on the LobbyMap.org website, which was not designed to be exhaustive. These limitations should be further investigated.
@@
4.8 Question Answering
@@
Table 10: Summary of datasets, labels, scopes, and sources for climate-related QA datasets.
@@
ClimaQA  [spokoyny2023answering] The text from a response to one of the CDP questions; and one of the questions from the CDP questionnaire 1: the response answers this question; 0: The response does not answer this question, but another one
ClimateQA [luccioni_analyzing_2020] Sentence from reports and a question based on the TCFD recommendations 1: the sentence answers the question; 0: The sentences does not answer the question
ClimaINS  [spokoyny2023answering] The text from a response to one of the questions from the NAIC questionnaire MANAGE, RISK PLAN, MITIGATE, ENGAGE, ASSESS, RISKS The labels correspond the 8 questions asked in the NAIC questionnaires
QA can be used for climate-specific applications such as structuring the information related to climate change from documents [luccioni_analyzing_2020, tobias_schimanski_climatebert-netzero_2023], building chatbots with climate-related knowledge to make it more accessible to non-experts [s_vaghefi_chatclimate_2023, cliamtebot_2022], or helping identify potentially misleading information [jingwei_ni_paradigm_2023].
@@
The QA process can be divided into two main steps [krausEnhancingLargeLanguage2023]: the retrieval step and the answer generation step. The retrieval step involves locating the relevant information or answer to the user’s query in external documents. The answer generation step involves formulating a response to the user’s query from (1) the information retrieved from documents in the first step or (2) the pre-trained internal knowledge of a model.
@@
Retriever systems can be generic, using techniques such as sentence similarity, BM25, or document tags to filter documents and retrieve specific passages. schimanski-etal-2024-climretrieve published ClimRetrieve, a benchmark for climate-related information retrieval in corporate reports. They found that simple embedding approaches are limited. Therefore, several works have specifically focused on improving the identification of answers within climate-related documents [luccioni_analyzing_2020, spokoyny2023answering]. luccioni_analyzing_2020 reformulated TCFD recommendations as questions and annotated reports to identify answers to the questions. The dataset includes questions paired with potential answers, each labeled to indicate whether the answer addresses the question or not. Based on the QA model trained on this dataset, they also proposed a tool called ClimateQA. spokoyny2023answering focused on questions from the CDP questionnaire and the NAIC Climate Risk Disclosure survey. As the responses to those questionnaires are publicly available, they constructed two datasets, ClimaINS and ClimaQA. ClimaQA is similar to ClimateQA (question/potential answer pairs). On the contrary, ClimaINS is a QA dataset framed as a classification task: it contains the responses from the NAIC survey, and each response is labeled as answering one of the 8 questions of the survey. Once a model has been trained on these datasets, it can be used to search for answers to the questions in other documents (effectively retrieving the information from an unstructured document). For ClimaINS, the authors experimented with multiple fine-tuned Transformers. As ClimaQA is framed as a retrieval task, the authors used BM25, sentence-BERT, and ClimateBERT to rank possible answers. Using the methodology evaluated on ClimaQA, they proposed a system to extract responses to CDP questions directly from unstructured reports and automatically fill the questionnaire [spokoyny2023answering]. luccioni_analyzing_2020 trained and evaluated a RoBERTa-based model called ClimateQA on finding answers to TCFD-based questions. They all reached good performance but with room for improvement,  [^7] See Table 21 in the Appendix for details. demonstrating the feasibility of the task but also the need for further research.
@@
After the retrieval, the second step of question answering is to generate an answer for the question (given input resources or not). The generation can be simple such as finding a particular information in a paragraph. For example, tobias_schimanski_climatebert-netzero_2023 used a QA model (RoBERTa SQuaD v2 [rajpurkar-etal-2016-squad]) to extract the target year, the percentage of reduction, and the baseline year from the reduction target of a company from an input text that contains an emission reduction target. While it is possible to rely on generalist models, as they are proficient few-shot learners [lm_few_shot_learner, thulke2024climategpt], multiple studies proposed using domain-specific models. Climate Bot [cliamtebot_2022] built a dataset of climate-specific QA. Given a question and a document (scientific/news), the model should find the span answering the question. mullappilly-etal-2023-arabic proposed a dataset of question-answer pairs based on ClimaBench [spokoyny2023answering] and CCMRC [cliamtebot_2022] to train a model to generate answers. thulke2024climategpt introduced a climate-specific corpus of prompt/completion pairs for Instruction Fine-Tuning created by experts and non-experts. cliamtebot_2022 finetuned an ALBERT model on their climate-specific QA dataset. mullappilly-etal-2023-arabic trained a Vicuna-7B on their climate-specific answer generation dataset. And thulke2024climategpt trained LLama-2-based models (ClimateGPT) on their domain-specific prompt/completion dataset. They also evaluated their ClimateGPT model alongside multiple LLMs on climate-related benchmarks. [^8] They found that the domain-specific models outperformed generalist ones (see Table 21 in the Appendix for details).
@@
By combining the last two components (a retriever and a QA system), one can develop Retrieval-Augmented Generation (RAG) systems. cliamtebot_2022 provide the first example of a RAG system in climate-related tasks based on sentence-BERT to retrieve documents and AlBERT to identify the answer. More recently, jingwei_ni_paradigm_2023 proposed ChatReport, a methodology based on ChatGPT for analyzing corporate reports through the lens of TCFD questions using RAG. The reports, the answers to the TCFD questions, and a TCFD conformity assessment are chunked and stored in a vector store for easy retrieval. Each chunk and/or answer can be retrieved and injected into the prompt. They added in the prompt the notions of greenwashing and cheap talk to invite the model to provide a critical analysis of the retrieved answers. Similarly, s_vaghefi_chatclimate_2023 introduced ChatClimate, a RAG-based pipeline to augment GPT-4 with knowledge about IPCC reports. While the previous LLM-based approaches rely on closed-source models, mullappilly-etal-2023-arabic and thulke2024climategpt proposed RAG-based pipelines relying on open-source models. cliamtebot_2022 is the only study that evaluated their RAG pipeline using classical metrics (F1-score, BLEU, METEOR). Unfortunately, those metrics are not well aligned with human judgment for text generation chen-etal-2019-evaluating. Therefore s_vaghefi_chatclimate_2023 and mullappilly-etal-2023-arabic relied on human or ChatGPT evaluations. They conclude that their approaches improve on non-domain-specific RAG systems.